In the Texas Lottery, six number are chosen between 1 and 54. to win the lottery, a person must match all sic number, but the order in which the number are chosen does NOT matter. Based on this information, compute the probability of winning the Texas lottery. Hint: The probability will be 1/(number of possible combinations of numbers)
Answer #1

since total number of ways to choose 6 number from 54 without concern of order =(54C6)

=54!/(6!*48!) =54*53*52*51*50*49/720=25827165

therefore  probability of winning the Texas lottery =1/25827165